Waterway, Inc., has bonds outstanding that will mature in 8 years. The bonds have a face value of $1,000. These bonds pay interest semiannually and have a coupon rate of 4.6 percent. If the bonds are currently selling at $901.92, what is the yield to maturity that an investor who buys them today can expect to earn? (Round answer to 1 decimal place, e.g. 5.2%.)

Yield to maturity Entry field with incorrect answer %

What is the effective annual yield? (Round answer to 3 decimal places, e.g. 5.275%.)

Effective annual yield Entry field with incorrect answer %
Answer #1

Given:

Face value of the Bond = $1000

Present Value of the bond = $901.92

Coupon Rate = 4.6% paid semi-annually

So, semi-annual rate = 2.3% ; Semi-annual coupon payment = 0.023*1000 = $23

Time to maturity = 8 years

Semi-annual period = 8*2 =16

Yield to maturity is calculated using excel RATE function

RATE(semiannual period, semi annual coupon payment, present value, face value,0)

RATE(16,23,-901.92,1000,0) =3.1%  

Semi-annual Yield to maturity = 3.1% or Annual Yield to Maturity - 3.1% *2 = 6.2%

Effective Annual Yield =[ (1 + semi-annual yield)^n - 1] where, n = compounding period

Effective Annual Yield = (1 + 0.031)^2 - 1 = 6.3%